QL=L^3+8L^2+12 then find optimal level of labour employment,optimal out level,maximum out put level&labour employment at which maximum out put is produced.MPL=0
1
Expert's answer
2018-02-12T07:25:08-0500
QL=-L^3+8L^2+12
(a) optimal level of labour employment is at:
QL' = 0,
-3L^2 + 16L = 0,
L(16 - 3L) = 0,
L = 0 or L = 5.33
(b) optimal output level can't be found without additional data.
(c) labour employment at which maximum out put is produced can't be found without additional data.